5. For this question, you will need a parent/guardian or a friend. Have this individual grab a handful of coins making sure there are only two types of coins in the group (i.e., nickels and dimes, quarters and pennies, pennies and dimes, etc). Your parent/guardian or friend should tell you the type of coins they’ve chosen, how many coins they have and the dollar amount of the group. From this information, you will set up two sets of equations and determine how many of each coin they have in their hand. Please send your instructor the name of the individual who helped you with this question, your two equations and the work you did to solve the system. (2 points)
Answer by richard1234(7193)

I am going to offer a theoretical example instead of actually doing it (I nor my friends would benefit or learn anything from doing it anyway).

Suppose your friend picks some quarters and pennies, and tells you that he has 11 coins worth 83 cents. If we let P and Q be the number of pennies and quarters, we set up a system of equations:

P + Q = 11 (since there are 11 coins)
P + 25Q = 83 (the value of the coins is 83 cents)

Here, we can subtract the first equation from the second equation to get (P + 25Q) - (P + Q) = 24Q = 83 - 11 = 72. Hence, 24Q = 72 --> Q = 3. We can replace Q = 3 into either equation (preferably 1st since it's easier) to solve for P. If we use the first equation, we get P + 3 = 11 --> P = 8, so three quarters and eight pennies.